### v2

# Locally similar distances and equality of the induced intrinsic distances

Abstract
Let $X$ be a set and $d_{1},d_{2}$ be two distances on $X$. We say that $d_{1}$ and $d_{2}$ are locally similar and write $d_{1}\cong d_{2}$ if $d_{1}$ and $d_{2}$ are topologically equivalent and, for every $a$ in $X$,
$\lim_{x\to a}\frac{d_{2}(x,a)}{d_{1}(x,a)}=1.$
We prove that if $d_{1}\cong d_{2}$, then the intrinsic distances induced by $d_{1}$ and $d_{2}$ coincide. We also provide sufficient conditions for $d_{1}\cong d_{2}$ and consider several examples related to reproducing kernel Hilbert spaces.
Keywords: intrinsic distance, equivalent distances, locally Lipschitz functions, distortion, local dilatation, reproducing kernel Hilbert spaces.
Mathematics Subject Classification (2020): 51K05, 30F45, 46E22.
1 Introduction
Background
The concept of distance is simple, but it plays a fundamental role in mathematics. See [11, 8, 12, 10] for many examples and applications of distances. There are many different notions of “equivalent distances” on a fixed set $X$, including the following standard notions:
  • $d_{1}$ and $d_{2}$ are topologically equivalent ($d_{1}\stackrel{{\scriptstyle\operatorname{top}}}{{\sim}}d_{2}$), if they induce the same topology;
  • $d_{1}$ and $d_{2}$ are uniformly equivalent ($d_{1}\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}d_{2}$), if they induce the same uniform structure, i.e., the identity function $(X,d_{1})\to(X,d_{2})$ and its inverse are uniformly continuous;
  • $d_{1}$ and $d_{2}$ are biLipschitz equivalent ($d_{1}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}$), if the identity function $(X,d_{1})\to(X,d_{2})$ and its inverse are Lipschitz continuous.
Every distance $d$ induces the intrinsic distance $d^{\ast}$, see [8, Chapter 2] or [12, Chapter 1]. Obviously, the binary relation $d_{1}\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}d_{2}$ (“intrinsic equivalence”), which we define by $d_{1}^{\ast}=d_{2}^{\ast}$, is an equivalence relation on the set of all distances defined on a fixed domain. A natural problem is to obtain criteria for $d_{1}^{\ast}=d_{2}^{\ast}$. It is easy to see that $d_{1}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}$ implies $d_{1}^{\ast}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}^{\ast}$ (Proposition 5.8), but it does not imply $d_{1}^{\ast}=d_{2}^{\ast}$ (Example 5.14).
Main results
In this paper, we introduce the concept of locally similar distances (Definition 2.1) and denote it by $d_{1}\cong d_{2}$. This concept is related to some concepts introduced by Capogna and Le Donne [9]; see Remark 2.5.
Our main result (Theorem 6.6) states that the relation $d_{1}\cong d_{2}$ is sufficient for $d_{1}^{\ast}=d_{2}^{\ast}$. We show that the relation $\cong$ is useful to compute $d^{\ast}$ in various examples. Nevertheless, $d_{1}\cong d_{2}$ is not necessary for $d_{1}^{\ast}=d_{2}^{\ast}$; see Examples 5.11 and 6.9.
By definition, locally similar distances are topologically equivalent. We show that $\cong$ is independent of $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$, and $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$: the biLipschitz equivalence does not imply the local similarity (Example 2.6), and the local similarity does not imply even the uniform equivalence (Example 2.8). Figure 1 shows the logical relationships among the equivalencies of distances considered in the paper.
$\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$ $\cong$ $\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{top}}}{{\sim}}$ Thm. 6.6Ex. 6.9Pr. 3.9Ex. 3.10Pr. 3.8Ex. 3.11Rem. 3.4(by def.)Ex. 2.6Ex. 2.8Ex. 5.12Ex. 5.14
Figure 1: Some logic relations between different equivalencies of distances. We give references to the corresponding theorems, propositions, remarks, and examples (counter-examples).
Several authors have studied the situation when $d$ is an intrinsic distance, i.e., $d^{\ast}=d$ (see [12, Section 1.3], [11, Chapter 1, parts A, B], [8, Section 2.3]). The constructions of this paper are useful in a more general case, namely when $d^{\ast}\cong d$.
We hope to relate this work to Riemannian metrics in future research. In this paper, we avoid Riemannian metrics to show the power of more elementary tools.
Structure of the paper
In Section 2, we introduce the concept of locally similar distances and show that it is not comparable with $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$ or $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$. In Section 3, we study two other concepts (strong local similarity $\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}$ and infinitesimal similarity $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$) that are slightly stronger than local similarity. In Section 4, we show that composing a distance $d_{1}$ with a function that satisfies certain properties yields a distance $d_{2}$ that is locally similar to $d_{1}$ ($d_{2}\cong d_{1}$, if $d_{2}=f\circ d_{1}$). In Section 5, we recall the definition of the intrinsic distance associated with a given distance; we also provide examples which are used throughout the paper. Section 6 contains a proof of the main result. Section 7 contains several examples related to the distances induced by kernels (or, equivalently, by reproducing kernel Hilbert spaces).
2 Locally similar distances
In this section, we introduce the main concept of this paper (Definition 2.1) and study its elementary properties. We have invented this concept as a sufficient condition for the intrinsic equivalence, see Theorem 6.6.
Definition 2.1.
Let $X$ be a set and $d_{1},d_{2}\colon X\to[0,+\infty)$ be distances on $X$. We say that $d_{1}$ and $d_{2}$ are locally similar and write $d_{1}\cong d_{2}$ if $d_{1}$ and $d_{2}$ are topologically equivalent and for every $a$ in $X$,
$\lim_{x\to a}\frac{d_{2}(x,a)}{d_{1}(x,a)}=1.$ (2.1)
In the context of Definition 2.1, let $\tau$ be the topology induced by $d_{1}$ (or $d_{2}$), and let $\tau(a)$ be the set of all open neighborhoods of $a$. Then, (2.1) can be written in the following more explicit form:
$\forall\varepsilon>0\qquad\exists U\in\tau(a)\qquad\forall x\in U\setminus\{a\}\qquad\left|\frac{d_{2}(x,a)}{d_{1}(x,a)}-1\right|<\varepsilon.$ (2.2)
Proposition 2.2.
Let $X$ be a set. The binary relation $\cong$, defined on the set of all distances on $X$, is an equivalence relation.
Proof.
It follows easily from the arithmetic properties of limits. ∎
Given a metric space $(X,d)$, for every $a$ in $X$ and every $\delta>0$ we denote by $B_{d}(a,\delta)$ the open ball
$B_{d}(a,\delta)\coloneqq\bigl\{x\in X\colon\ d(x,a)<\delta\bigr\}.$
Furthermore, we denote by $\tau_{d}$ the topology induced by $d$.
In the next proposition, we characterize the relation $d_{1}\cong d_{2}$ avoiding the induced topology.
Proposition 2.3.
Let $X$ be a set and $d_{1},d_{2}\colon X\to[0,+\infty)$ be distances on $X$. $d_{1}\cong d_{2}$ if and only if for every $a$ in $X$, the following limit relations hold:
$\lim_{\begin{subarray}{c}d_{1}(x,a)\to 0\\ x\neq a\end{subarray}}\frac{d_{2}(x,a)}{d_{1}(x,a)}=1,$ (2.3)
$\lim_{\begin{subarray}{c}d_{2}(x,a)\to 0\\ x\neq a\end{subarray}}\frac{d_{1}(x,a)}{d_{2}(x,a)}=1.$ (2.4)
We understand (2.3) and (2.4) in the following sense, respectively:
$\forall\varepsilon>0\qquad\exists\delta>0\qquad\forall x\in B_{d_{1}}(a,\delta)\setminus\{a\}\qquad\left|\frac{d_{2}(x,a)}{d_{1}(x,a)}-1\right|<\varepsilon,$ (2.5)
$\forall\varepsilon>0\qquad\exists\delta>0\qquad\forall x\in B_{d_{2}}(a,\delta)\setminus\{a\}\qquad\left|\frac{d_{1}(x,a)}{d_{2}(x,a)}-1\right|<\varepsilon.$ (2.6)
Proof.
Necessity. Suppose that $d_{1}\cong d_{2}$, i.e., (2.1) holds for each $a$. Since $\tau_{d_{1}}$ is induced by $d_{1}$, there exists $\delta>0$ such that $B_{d_{1}}(a,\delta)\subseteq U$. Thereby, we get (2.5). By the arithmetic properties of limits, (2.1) implies a similar limit with the quotient $d_{1}/d_{2}$, which yields (2.6).
Sufficiency. Suppose that (2.5) and (2.6) are fulfilled for every $a$ in $X$. Let $a\in X$. Applying (2.5) with $\varepsilon=1/2$ we get $\delta>0$ such that for all $x$ in $B_{d_{1}}(a,\delta)$,
$d_{2}(x,a)\leq\frac{3}{2}d_{1}(x,a),$
which implies that $B_{d_{1}}(a,r)\subseteq B_{d_{2}}(a,3r/2)$ for every $r$ with $0<r<\delta$. Since $a$ is arbitrary, we get $\tau_{d_{2}}\subseteq\tau_{d_{1}}$. In a similar way, using (2.4), we obtain that $\tau_{d_{1}}\subseteq\tau_{d_{2}}$. So, $d_{1}$ and $d_{2}$ induce the same topology. Finally, taking $U\coloneqq B_{d_{1}}(a,\delta)$, we see that (2.5) implies (2.2). ∎
Remark 2.4.
This paper could be easily generalized to pseudodistances. Instead of $x\neq a$, we should require $d_{1}(x,a)>0$ in (2.1), (2.2), (2.3), and (2.5), and $d_{2}(x,a)>0$ in (2.4) and (2.6). Of course, the topological equivalence of $d_{1}$ and $d_{2}$ implies that $d_{1}(x,a)>0$ if and only if $d_{2}(x,a)>0$.
Remark 2.5.
The binary relation $\cong$ can be described in terms of some concepts from Capogna and Le Donne [9, Subsection 2.1]. Given two metric spaces $(X,d_{1})$ and $(Y,d_{2})$, a homeomorphism $f\colon X\to Y$, and a point $a$ in $X$, they consider the following superior and inferior limits:
$L_{f}(a)\coloneqq\limsup_{x\to a}\frac{d_{2}(f(x),f(a))}{d_{1}(x,a)},\qquad\ell_{f}(a)\coloneqq\liminf_{x\to a}\frac{d_{2}(f(x),f(a))}{d_{1}(x,a)}.$ (2.7)
We apply these concepts in the particular case where $Y=X$ and $f$ is the identity function $\operatorname{id}_{X}\colon X\to X$ where the domain is equipped with $d_{1}$ and the codomain is equipped with $d_{2}$. It is well known that the existence of $\lim$ is equivalent to the equality between $\limsup$ and $\liminf$; see, e.g., [7, Chapter IV, Subsection 5.6]. Thus, $d_{1}\cong d_{2}$ if and only if for every $a$ in $X$, $L_{\operatorname{id}_{X}}(a)=\ell_{\operatorname{id}_{X}}(a)=1$. In the terminology from [9], this means that $\operatorname{id}_{X}$ has pointwise Lipschitz constant $1$ and distortion $1$ at every point.
Example 2.6.
There exist two distances that are biLipschitz equivalent, but not locally similar. Consider $X=[0,1]$. Define $\rho\colon X^{2}\to[0,+\infty)$, $\rho(x,y)\coloneqq|f(x)-f(y)|$, where
$f(x)\coloneqq x(1+x).$
More explicitly,
$\rho(x,y)=|x-y|(1+x+y).$
Let $d_{[0,1]}$ be the Euclidean distance on $[0,1]$. Obviously, $\rho$ is biLipschitz equivalent to $d_{[0,1]}$:
$|x-y|\leq\rho(x,y)\leq 3|x-y|.$
We observe that the limit from Definition 2.1 exists for each $a$ in $X$, but it depends on $a$:
$\lim_{x\to a}\frac{\rho(x,a)}{d_{[0,1]}(x,a)}=\lim_{x\to a}(1+x+a)=1+2a.$
Therefore, $\rho$ is not locally similar to $d_{[0,1]}$. Moreover, there is no $C>0$ such that $\rho\cong Cd_{[0,1]}$. ∎
Example 2.7.
There exist two distances that are locally similar, but not biLipschitz equivalent. Consider $(\mathbb{R},d_{\mathbb{R}})$ and $(\mathbb{R},\rho)$, where
$\rho(x,y)\coloneqq\min\{1,|x-y|\}.$
Then, $d_{\mathbb{R}}\cong\rho$ but there is no $C>0$ such that $d_{\mathbb{R}}(x,y)\leq C\rho(x,y)$ for every $x,y$ in $\mathbb{R}$. ∎
Example 2.8.
There exist two distances that are locally similar, but not uniformly equivalent. We consider the open interval $X=(0,2\pi)$ provided with the Euclidean distance $d_{X}(a,b)\coloneqq|a-b|$ and the “circular distance”
$\rho(a,b)\coloneqq\bigl|\operatorname{e}^{\operatorname{i}a}-\operatorname{e}^{\operatorname{i}b}\bigr|,\qquad\text{i.e.},\qquad\rho(a,b)=2\left|\sin\frac{a-b}{2}\right|,$
see Figure 2. It is easy to see that $d_{X}\cong\rho$. Nevertheless, for $x_{m}=\frac{1}{m}$ and $y_{m}=2\pi-\frac{1}{m}$,
$\lim_{m\to\infty}\rho(x_{m},y_{m})=0,\qquad\lim_{m\to\infty}d_{X}(x_{m},y_{m})=2\pi.$
Therefore, the identity function $(X,\rho)\to(X,d_{X})$ is not uniformly continuous. Hence, $d_{X}$ and $\rho$ are not uniformly equivalent. ∎
[…]
Figure 2: Distances from Example 2.8.
3 Strongly locally similar and infinitesimally similar distances
In this section, we give two alternative concepts for local similarity of distances (Definitions 3.1 and 3.6) and show that they are slightly stronger than $\cong$ from Definition 2.1. The concepts introduced in this section are close to definitions from Gromov [11, Definition 1.1] and Lytchak [13, Definition 3.2].
Strongly locally similar distances
Definition 3.1.
Let $X$ be a set and $d_{1},d_{2}\colon X\to[0,+\infty)$ be distances on $X$. We say that $d_{1}$ and $d_{2}$ are strongly locally similar and write $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$ if $d_{1}$ and $d_{2}$ are topologically equivalent and for every $a$ in $X$,
$\lim_{\begin{subarray}{c}(x,y)\to(a,a)\\ x\neq y\end{subarray}}\frac{d_{2}(x,y)}{d_{1}(x,y)}=1.$ (3.1)
In the context of Definition 3.1, let $\tau$ be the topology induced by $d_{1}$ (or $d_{2})$, and let $\tau(a)$ be the set of all open neighborhoods of $a$. Furthermore, let $\Delta(X)$ be the “diagonal” of $X^{2}$:
$\Delta(X)\coloneqq\{(x,y)\in X^{2}\colon\ x=y\}.$
The limit in (3.1) is understood in the sense of the product topology on $X^{2}$. More explicitly, (3.1) means that
$\forall\varepsilon>0\qquad\exists U\in\tau(a)\qquad\forall(x,y)\in U^{2}\setminus\Delta(X)\qquad\left|\frac{d_{2}(x,y)}{d_{1}(x,y)}-1\right|<\varepsilon.$ (3.2)
Proposition 3.2.
Let $X$ be a set. The binary relation $\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}$, defined on the set of all distances on $X$, is an equivalence relation.
Proof.
It follows easily from the arithmetic properties of limits. ∎
In the next proposition, we characterize the condition $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$ avoiding the induced topology.
Proposition 3.3.
Let $X$ be a set and $d_{1},d_{2}\colon X\to[0,+\infty)$ be distances on $X$. Then, $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$ if and only if for every $a$ in $X$, the following limit relations hold:
$\lim_{\begin{subarray}{c}d_{1}(x,a)\to 0\\ d_{1}(y,a)\to 0\\ x\neq y\end{subarray}}\frac{d_{2}(x,y)}{d_{1}(x,y)}=1,$ (3.3)
$\lim_{\begin{subarray}{c}d_{2}(x,a)\to 0\\ d_{2}(y,a)\to 0\\ x\neq y\end{subarray}}\frac{d_{1}(x,y)}{d_{2}(x,y)}=1.$ (3.4)
We understand (3.3) and (3.4) in the following sense, respectively:
$\forall\varepsilon>0\qquad\exists\delta>0\qquad\forall(x,y)\in B_{d_{1}}(a,\delta)^{2}\setminus\Delta(X)\qquad\left|\frac{d_{2}(x,y)}{d_{1}(x,y)}-1\right|<\varepsilon,$ (3.5)
$\forall\varepsilon>0\qquad\exists\delta>0\qquad\forall(x,y)\in B_{d_{2}}(a,\delta)^{2}\setminus\Delta(X)\qquad\left|\frac{d_{1}(x,y)}{d_{2}(x,y)}-1\right|<\varepsilon.$ (3.6)
Proof.
Necessity. Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$, i.e., (3.2) holds for every $a$ in $X$. Since $\tau_{d_{1}}$ is induced by $d_{1}$, there exists $\delta>0$ such that $B_{d_{1}}(a,\delta)\subseteq U$. Thereby we get (3.5). By the arithmetic properties of limits, (3.1) implies a similar limit with the quotient $d_{1}/d_{2}$, which yields (3.6).
Sufficiency. Suppose that (3.5) and (3.6) are fulfilled for every $a$ in $X$. Applying (3.5) with $\varepsilon=1/2$ and $y=a$ we get $\delta>0$ such that for all $x$ in $B_{d_{1}}(a,\delta)$,
$d_{2}(x,a)\leq\frac{3}{2}d_{1}(x,a),$
which implies that $B_{d_{1}}(a,r)\subseteq B_{d_{2}}(a,3r/2)$ for every $r$ with $0<r<\delta$. Since $a$ is arbitrary, we get $\tau_{d_{2}}\subseteq\tau_{d_{1}}$. In a similar way, using (3.6),we obtain that $\tau_{d_{1}}\subseteq\tau_{d_{2}}$. So, $d_{1}$ and $d_{2}$ induce the same topology. Finally, taking $U\coloneqq B_{d_{1}}(a,\delta)$ we see that (3.5) implies (3.2). ∎
Remark 3.4.
In Example 2.8, the distances $d_{X}$ and $\rho$ are strongly locally similar but not uniformly equivalent.
Description of strong local similarity in terms of local dilatation
Suppose that $(X,d_{X})$ and $(Y,d_{Y})$ are metric spaces, $f\colon X\to Y$, and $a\in X$. According to Gromov [11, Definition 1.1] the local dilatation of $f$ at the point $a$ is defined as
$\operatorname{dil}_{a}(f)\coloneqq\lim_{\delta\to 0}\;\sup_{\begin{subarray}{c}x,y\in B_{d_{1}}(a,\delta)\\ x\neq y\end{subarray}}\frac{d_{2}(f(x),f(y))}{d_{1}(x,y)}.$ (3.7)
This expression can be written as the following upper limit:
$\operatorname{dil}_{a}(f)=\limsup_{\begin{subarray}{c}d_{1}(x,a)\to 0\\ d_{1}(y,a)\to 0\\ (x,y)\notin\Delta(X)\end{subarray}}\frac{d_{2}(f(x),f(y))}{d_{1}(x,y)}.$ (3.8)
Proposition 3.5.
Let $X$ be a set and $d_{1},d_{2}$ be distances on $X$. Then, the following conditions are equivalent:
  • (a)
    $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$;
  • (b)
    the identity function $(X,d_{1})\to(X,d_{2})$ and its inverse $(X,d_{2})\to(X,d_{1})$ have local dilatation $1$ at every point of $X$.
Proof.
[…]
$\lim_{\delta\to 0}\ \sup_{\begin{subarray}{c}x,y\in B_{d_{1}}(a,\delta)\\ x\neq y\end{subarray}}\frac{d_{2}(x,y)}{d_{1}(x,y)}=1$ (3.9)
and
$\lim_{\delta\to 0}\ \sup_{\begin{subarray}{c}x,y\in B_{d_{2}}(a,\delta)\\ x\neq y\end{subarray}}\frac{d_{1}(x,y)}{d_{2}(x,y)}=1.$ (3.10)
For every $a$ in $X$, the limit relations (3.9) and (3.10) imply that for every $\varepsilon>0$ there exists $\delta>0$ such that
$\forall(x,y)\in B_{d_{1}}(a,\delta)^{2}\setminus\Delta(X)\qquad\frac{d_{2}(x,y)}{d_{1}(x,y)}<1+\varepsilon$ (3.11)
and
$\forall(x,y)\in B_{d_{2}}(a,\delta)^{2}\setminus\Delta(X)\qquad\frac{d_{1}(x,y)}{d_{2}(x,y)}<1+\varepsilon.$ (3.12)
Similarly to the proof of Proposition 3.3, we can deduce that $\tau_{d_{1}}=\tau_{d_{2}}$. After that, for each $a$ in $X$, from (3.11) and (3.12) we get $U\coloneqq B_{d_{1}}(a,\delta)\cap B_{d_{2}}(a,\delta)$ in $\tau(a)$ such that
$\forall(x,y)\in U^{2}\setminus\Delta(X)\qquad\frac{1}{1+\varepsilon}<\frac{d_{2}(x,y)}{d_{1}(x,y)}<1+\varepsilon.$
Therefore, (3.2) holds.
[…]
Infinitesimally similar distances
The following concept is inspired by Lytchak [13, Definition 3.2].
Definition 3.6.
Let $X$ be a set and $d_{1},d_{2}\colon X\to[0,+\infty)$ be distances on $X$. We say that $d_{1}$ and $d_{2}$ are infinitesimally similar and write $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$ if $d_{1}$ and $d_{2}$ are topologically equivalent and for every $a$ in $X$,
$\lim_{(x,y)\to(a,a)}\frac{|d_{2}(x,y)-d_{1}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}=0.$ (3.13)
As usual, in the definition of $\lim_{(x,y)\to(a,a)}$ we assume that $(x,y)\neq(a,a)$. To include the point $(a,a)$, we rewrite the limit relation (3.13) in the following equivalent form:
$\forall\varepsilon>0\quad\exists V\in\tau(a)\quad\forall x,y\in V\qquad|d_{2}(x,y)-d_{1}(x,y)|\leq\varepsilon\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr),$ (3.14)
where $\tau$ is the topology induced by $d_{1}$ (or by the topologically equivalent distance $d_{2}$).
Proposition 3.7.
Let $X$ be a set. The binary relation $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$, defined on the set of all distances on $X$, is an equivalence relation.
Proof.
Obviously, $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$ is reflexive.
1. Symmetry. Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$. Let $\tau\coloneqq\tau_{d_{1}}$; by definition of $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$, we also have $\tau=\tau_{d_{2}}$. Let $a\in X$ and $\varepsilon>0$. From (3.14), there is $V$ in $\tau(a)$ such that for every $x,y$ in $V$,
$|d_{1}(x,y)-d_{2}(x,y)|\leq\frac{\min\{\varepsilon,1\}}{2}\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr).$ (3.15)
Taking in (3.15) $y=a$ and then $x=a$, we get
$d_{1}(x,a)-d_{2}(x,a)\leq\frac{d_{1}(x,a)}{2},\qquad d_{1}(y,a)-d_{2}(y,a)\leq\frac{d_{1}(y,a)}{2}.$
We sum the last two inequalities and simplify:
$d_{1}(x,a)+d_{1}(y,a)\leq 2\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr).$ (3.16)
Thus, for every $x,y$ in $V$, by (3.15) and (3.16),
$|d_{2}(x,y)-d_{1}(x,y)|\leq\frac{\varepsilon}{2}\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr)\leq\varepsilon\,\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr).$
2. Transitivity. Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$ and $d_{2}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{3}$. Let $\tau\coloneqq\tau_{d_{1}}$. Then, $\tau=\tau_{d_{2}}=\tau_{d_{3}}$. By symmetry, we have $d_{2}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{1}$. Let $a\in X$ and $\varepsilon>0$. There exists $V$ in $\tau(a)$ such that for all $x,y,z$ in $V$,
$\bigl|d_{1}(x,y)-d_{2}(x,y)\bigr|$ $\leq\frac{\varepsilon}{3}\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr),$
$\bigl|d_{2}(x,y)-d_{1}(x,y)\bigr|$ $\leq\frac{1}{2}\,\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr),$
and
$\bigl|d_{2}(x,y)-d_{3}(x,y)\bigr|\leq\frac{\varepsilon}{3}\,\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr).$
The second of these inequalities implies an analog of (3.16), with swapped $d_{1}$ and $d_{2}$:
$d_{2}(x,a)+d_{2}(y,a)\leq 2\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr).$
Thus,
$\bigl|d_{1}(x,y)-d_{3}(x,y)\bigr|$ $\leq\bigl|d_{1}(x,y)-d_{2}(x,y)\bigr|+\bigl|d_{2}(x,y)-d_{3}(x,y)\bigr|$
$\leq\frac{\varepsilon}{3}\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr)+\frac{\varepsilon}{3}\,\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr)$
$\leq\varepsilon\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr).\qed$
Comparison of strong local similarity, infinitesimal similarity, and local similarity
Proposition 3.8.
Let $X$ be a set and $d_{1},d_{2}$ be distances on $X$. If $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$, then $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$.
Proof.
Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$. Let $a\in X$. We use the following upper bound for the expression in (3.13):
$\frac{|d_{2}(x,y)-d_{1}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}\leq\frac{|d_{2}(x,y)-d_{1}(x,y)|}{d_{1}(x,y)}=\left|\frac{d_{2}(x,y)}{d_{1}(x,y)}-1\right|.$
As $(x,y)\to(a,a)$ and $x\neq y$, the last expression tends to zero. ∎
Proposition 3.9.
Let $X$ be a set and $d_{1},d_{2}$ be distances on $X$. If $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$, then $d_{1}\cong d_{2}$.
Proof.
Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$. Let $a\in X$. We suppose that the limit relation (3.13) holds. Fixing $y=a$ we get the following consequence of (3.13):
$\lim_{x\to a}\frac{|d_{2}(x,a)-d_{1}(x,a)|}{d_{1}(x,a)}=0.$
It is equivalent to (2.1). ∎
The next example shows, in particular, that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$ is not equivalent to $d_{1}\cong d_{2}$.
Example 3.10.
Let $X$ be the following “hook” in $\mathbb{R}^{2}$:
$X=\Bigl([0,+\infty)\times\{0\}\Bigr)\cup\Bigl(\{0\}\times[0,+\infty)\Bigr)=\{(x,0)\colon\ x\geq 0\}\cup\{(0,y)\colon\ y\geq 0\}.$
We denote by $d_{1}$ the “taxi distance” on $X$ and by $d_{2}$ the Euclidean distance on $X$:
$d_{1}((u,v),(r,s))=|u-r|+|v-s|,\qquad d_{2}((u,v),(r,s))=\sqrt{(u-r)^{2}+(v-s)^{2}}.$
For two points belonging to the same halfaxis, we have
$d_{1}((u,0),(r,0))=|u-r|=d_{2}((u,0),(r,0)),$
while for two points belonging to the different halfaxes,
$d_{1}((u,0),(0,s))=u+s,\qquad d_{2}((u,0),(0,s))=\sqrt{u^{2}+s^{2}}.$
See Figure 3.
Notice that if $a=(0,0)$, then $d_{2}(x,a)=d_{1}(x,a)$ for every $x$ in $X$. If $a\neq(0,0)$, then $d_{2}(x,a)=d_{1}(x,a)$ for every $x$ sufficiently close to $a$. Therefore, $d_{1}\cong d_{2}$.
It is also easy to see that $d_{1}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}$ and $d_{2}^{\ast}=d_{1}^{\ast}=d_{1}$.
On the other hand, working with $a=(0,0)$, $x=(u,0)$, and $y=(0,u)$, we get for every $u>0$:
$\frac{d_{2}((u,0),(0,u))}{d_{1}((u,0),(0,u))}=\frac{\sqrt{2}\,u}{2u}=\frac{\sqrt{2}}{2}$
and
$\frac{|d_{2}((u,0),(0,u))-d_{1}((u,0),(0,u))|}{d_{1}((u,0),(0,0))+d_{1}((0,u),(0,0))}=\frac{2-\sqrt{2}}{2}.$
Therefore, $d_{1}$ and $d_{2}$ are not strongly locally similar nor infinitesimally similar. ∎
a=(0,0)a=(0,0)x=(u,0)x=(u,0)$y=(0,s)$
Figure 3: Distances from Example 3.10.
Omer Cantor (https://orcid.org/0009-0007-2969-7170) constructed the following example and explained it to us. It shows that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$ is not equivalent to $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$.
Example 3.11.
For each parameter $\alpha$ satisfying $0<\alpha\leq 1$, we define $\rho_{\alpha}\colon[-1,1]^{2}\to[0,+\infty)$ by
$\rho_{\alpha}(x,y)\coloneqq\begin{cases}|x-y|,&xy\geq 0;\\ |x+y|+\alpha\bigl(\min\{|x|,|y|\}\bigr)^{2},&xy<0.\end{cases}$
Intuitively, the definition of $\rho_{\alpha}$ means that for each $u$ in $(0,1]$ there is a “teleportation” of cost $\alpha u^{2}$ between $u$ and u-u. Thus, if $0<u<v\leq 1$, then $\rho_{\alpha}(-u,v)$ as obtained as the sum of two terms: the teleportation between u-u and $u$ and the usual walk between $u$ and $v$. It is easy to verify that $\rho_{\alpha}$ is a distance on [1,1][-1,1]. It induces the usual topology on [1,1][-1,1]. Moreover, if $a\neq 0$, then there exists an open neighborhood $V$ of $a$ such that $\rho_{\alpha}(x,y)=|x-y|$ for each $x,y$ in $V$. Consider two different distances from this family, say $\rho_{1/2}$ and $\rho_{1}$. For $a=0$ and $x,y\in[-1,1]$ such that $xy<0$, we have
$\frac{|\rho_{1/2}(x,y)-\rho_{1}(x,y)|}{\rho_{1}(x,0)+\rho_{1}(0,y)}=\frac{\frac{1}{2}\bigl(\min\{|x|,|y|\}\bigr)^{2}}{|x|+|y|}.$
Obviously, this expression tends to 0 as $(x,y)\to(0,0)$. For all other choices of $a,x,y$, the corresponding quotient is equal to 0. Therefore, $\rho_{1/2}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}\rho_{1}$.
On the other hand, for $a=0$, $x_{k}=1/k$, and $y_{k}=-1/k$, we consider the quotient
$\frac{\rho_{1/2}(x_{k},y_{k})}{\rho_{1}(x_{k},y_{k})}=\frac{\frac{1}{2}(1/k)^{2}}{(1/k)^{2}}=\frac{1}{2}.$
As $k\to\infty$, we have $(x_{k},y_{k})\to(0,0)$, but the quotient above does not tend to $1$. Hence, $\rho_{1/2}$ and $\rho_{1}$ are not strongly locally similar.
4 Locally similar distances and composition
In this section, we consider a particular situation where two distances $d_{1}$ and $d_{2}$ are related by $d_{2}=f\circ d_{1}$. We provide a sufficient condition for $d_{2}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{1}$ in terms of $f$. More precisely, for convenience in applications, we provide a sufficient condition for $d_{2}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}Qd_{1}$, with $0<Q<+\infty$. According to Propositions 3.8 and 3.9, our condition is also sufficient for $d_{2}\cong Qd_{1}$.
Lemma 4.1.
Let $f\colon[0,+\infty)\to[0,+\infty)$ and $C>0$ be such that $f(t)\geq C\min\{t,1\}$ for every $t\geq 0$. Then
  • 1)
    for every $t>0$, $f(t)>0$;
  • 2)
    $\lim_{f(t)\to 0^{+}}t=0$, i.e.,
    $\forall\eta>0\quad\exists\xi>0\quad\forall t\in[0,+\infty)\quad\Bigl(0<f(t)<\xi\quad\Longrightarrow\quad 0<t<\eta\Bigr).$
Proof.
Property 1) follows directly from the hypothesis. We are going to prove 2). Let $\eta>0$. Define $\xi\coloneqq\min\{C\eta,C/2\}$. Let $t$ be a positive number such that $0<f(t)<\xi$. Notice that $t<1$, since otherwise we get:
$C=C\min\{t,1\}\leq f(t)<\frac{C}{2}.$
Therefore, the result follows:
$t=\min\{t,1\}\leq\frac{f(t)}{C}<\frac{\xi}{C}\leq\eta.\qed$
Proposition 4.2.
Let $(X,d_{1})$ be a metric space, $0<Q<+\infty$, and $f\colon[0,+\infty)\to[0,+\infty)$ be a function with the following properties:
  1. 1)
    $f(0)=0$;
  2. 2)
    $\lim_{t\to 0^{+}}\frac{f(t)}{t}=Q$;
  3. 3)
    there is a constant $C>0$ such that for every $t\geq 0$, $f(t)\geq C\min\{t,1\}$.
Define $d_{2}\coloneqq f\circ d_{1}$. Suppose that $d_{2}$ is a distance. Then, $d_{2}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}Qd_{1}$ and $d_{2}\cong Qd_{1}$.
Proof.
Part 1. Let us verify an analog of (3.3) for $Qd_{1}$ and $d_{2}$. Given $\varepsilon>0$, by 2), there is $\delta_{1}>0$ such that, if $0<t<\delta_{1}$, then
$\left|\frac{f(t)}{t}-Q\right|<Q\varepsilon.$
Define $\delta_{2}\coloneqq\frac{Q\delta_{1}}{2}$. Let $a\in X$ and $x,y\in X$ such that
$x\neq y,\qquad Qd_{1}(x,a)<\delta_{2},\qquad Qd_{1}(y,a)<\delta_{2}.$
Put $s\coloneqq d_{1}(x,y)$. Then $0<s\leq d_{1}(x,a)+d_{1}(a,y)<2\delta_{2}/Q=\delta_{1}$ and
$\left|\frac{d_{2}(x,y)}{Qd_{1}(x,y)}-1\right|=\frac{1}{Q}\left|\frac{d_{2}(x,y)}{d_{1}(x,y)}-Q\right|=\frac{1}{Q}\left|\frac{f(s)}{s}-Q\right|<\varepsilon.$
Part 2. Let us verify an analog of (3.4) for $Qd_{1}$ and $d_{2}$. By property 2),
$\lim_{t\to 0^{+}}\frac{t}{f(t)}=\frac{1}{Q}.$
Given $\varepsilon>0$, there is $\delta_{3}>0$ such that if $0<t<\delta_{3}$, then
$\left|\frac{t}{f(t)}-\frac{1}{Q}\right|<\frac{\varepsilon}{Q}.$
From Proposition 4.1, there is $\delta_{4}>0$ such that if $0<f(t)<\delta_{4}$, then $0<t<\delta_{3}$.
Let $a\in X$ and $x,y\in X$ such that $x\neq y$, $d_{2}(x,a)<\delta_{4}/2$, and $d_{2}(y,a)<\delta_{4}/2$. Put $s\coloneqq d_{1}(x,y)$. Then $0<f(s)=d_{2}(x,y)\leq d_{2}(x,a)+d_{2}(a,y)<\delta_{4}$. Therefore, $0<s<\delta_{3}$, and
$\left|\frac{Qd_{1}(x,y)}{d_{2}(x,y)}-1\right|=Q\left|\frac{d_{1}(x,y)}{d_{2}(x,y)}-\frac{1}{Q}\right|=Q\left|\frac{s}{f(s)}-\frac{1}{Q}\right|<\varepsilon.\qed$
Corollary 4.3.
Let $(X,d_{1})$ be a metric space, $0<Q<+\infty$, and $f\colon[0,+\infty)\to[0,+\infty)$ be a concave increasing function such that $f(0)=0$ and
$\lim_{t\to 0^{+}}\frac{f(t)}{t}=Q.$
Define $d_{2}\coloneqq f\circ d_{1}$. Then, $d_{2}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}Qd_{1}$ and $d_{2}\cong Qd_{1}$.
Proof.
The assumptions on $f$ imply that $f(t)>0$ for $t>0$. Moreover, $d_{2}$ is a distance because $f$ is concave and increasing, and $f(0)=0$ (see a similar fact in [10, Section 4.1]). We will verify property 3) from Proposition 4.2. Let $t\geq 0$. If $t<1$, then
$tf(1)=(1-t)f(0)+tf(1)\leq f(t).$
If $t\geq 1$, then $f(1)\leq f(t)$, because $f$ is increasing. Joining these two cases, for every $t\geq 0$ we get
$f(1)\min\{t,1\}\leq f(t).\qed$
5 The intrinsic distance induced by a given distance
In this section, we recall the definition of the intrinsic distance $d^{\ast}$ induced by a given distance $d$. It is a standard construction in the theory of length spaces (see, for example, Gromov [11, Example 1.4 (a)], Burago, Burago, and Ivanov [8, Section 2.3], Kunzinger and Steinbauer [12, Section 1.2.3]). In this section, we also compute $d^{\ast}$ for several examples. In particular, we show that $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$ does not imply $\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}$ and $\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}$ does not imply $\stackrel{{\scriptstyle\operatorname{top}}}{{\sim}}$.
We now fix some notation for the intermediate concepts which will be used in the examples and the proof of the main result.
Definition 5.1 (continuous paths between two points).
Let $(X,d)$ be a metric space. For each $x,y$ in $X$, define
$\Gamma_{d}(x,y)\coloneqq\bigl\{\gamma\in C([0,1],X,d)\colon\quad\gamma(0)=x,\quad\gamma(1)=y\bigr\}.$
Here $C([0,1],X,d)$ stands for the set of all functions $\gamma\colon[0,1]\to X$ that are continuous with respect to $\tau_{d}$.
Definition 5.2 (Riemannian partitions of an interval).
Let $a,b]\subseteq\mathbb{R}$ be an arbitrary interval. We denote by $\mathcal{P}([a,b])$ the set of all finite subsets of $[a,b]$ that include $a$ and $b$. In particular, we denote $\mathcal{P}([0,1])$ by $\mathcal{P}$.
Given $P$ in $\mathcal{P}([a,b])$, we enumerate its elements in ascending order:
$P=\{t_{0},t_{1},\ldots,t_{m}\},\qquad a=t_{0}<t_{1}<\cdots<t_{m}=b.$ (5.1)
Definition 5.3 (the polygonal length of a path with respect to a partition).
Let $(X,d)$ be a metric space, $x,y\in X$, $\gamma\in\Gamma_{d}(x,y)$, and $P\in\mathcal{P}$ be of the form (5.1). Define
$S(d,\gamma,P)\coloneqq\sum_{k=0}^{m-1}d(\gamma(t_{k}),\gamma(t_{k+1})).$
Lemma 5.4.
Let $(X,d)$ be a metric space, $x,y\in X$, and $\gamma\in\Gamma_{d}(x,y)$. If $P,Q\in\mathcal{P}$ such that $P\subseteq Q$, then
$S(d,\gamma,P)\leq S(d,\gamma,Q).$
Proof.
It follows by an iterative application of the triangle inequality. ∎
Definition 5.5 (the length of a path with respect to a distance).
Let $(X,d)$ be a metric space, $x,y\in X$, and $\gamma\in\Gamma_{d}(x,y)$. Define
$L(d,\gamma)\coloneqq\sup_{P\in\mathcal{P}}S(d,\gamma,P).$
Definition 5.6 (the intrinsic distance induced by a distance).
Let $(X,d)$ be a metric space. Define $d^{\ast}\colon X^{2}\to[0,+\infty]$ by the following rule:
$d^{\ast}(x,y)\coloneqq\inf_{\gamma\in\Gamma_{d}(x,y)}L(d,\gamma).$
Following Definition 5.6, if $\Gamma_{d}(x,y)=\emptyset$ (i.e., if there are no continuous paths between $x$ and $y$), then $d^{\ast}(x,y)=+\infty$.
It is well known that $d^{\ast}$ is a $0,+\infty]-$valued distance on $X$, such that $d^{\ast}(x,y)\geq d(x,y)$ for every $x,y$ in $X$. Moreover, it is well known that $(d^{\ast})^{\ast}=d^{\ast}$. These facts are proved in the books cited at the beginning of this section.
Remark 5.7.
In this paper, we define paths as continuous functions with domain $[0,1]$. Similarly, one can consider paths defined on arbitrary closed intervals of the form $[a,b]$. Such generalization does not change $L(d,\gamma)$ or $d^{\ast}(x,y)$, since any path defined on an arbitrary interval $[a,b]$ can be reparametrized to $[0,1]$.
Proposition 5.8.
Let $d_{1},d_{2}$ be distances on $X$ such that $d_{1}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}$. Then $d_{1}^{\ast}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}^{\ast}$.
Proof.
The proof is a simple exercise on supremums and infimums; we give it for the sake of completeness. By the definition of $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$, there exist $C_{1},C_{2}$ in $(0,+\infty)$ such that $d_{1}\leq C_{1}d_{2}$ and $d_{2}\leq C_{2}d_{1}$. Notice that $\Gamma_{d_{1}}(x,y)=\Gamma_{d_{2}}(x,y)$ for every $x,y$ in $X$, because $d_{1}\stackrel{{\scriptstyle\operatorname{top}}}{{\sim}}d_{2}$.
Let $x,y\in X$ and $\gamma\in\Gamma_{d_{1}}(x,y)$. For every $P$ in $\mathcal{P}$,
$S(d_{2},\gamma,P)\leq C_{2}S(d_{1},\gamma,P)\leq C_{2}L(d_{1},\gamma).$
Passing to the supremum over $P$ we conclude that $L(d_{2},\gamma)\leq C_{2}L(d_{1},\gamma)$. So, for every $\gamma$ in $\Gamma_{d_{1}}(x,y)$,
$\frac{1}{C_{2}}d_{2}^{\ast}(x,y)\leq\frac{1}{C_{2}}L(d_{2},\gamma)\leq L(d_{1},\gamma).$
Passing to the infimum over $\gamma$ in $\Gamma_{d_{1}}(x,y)$ we get $d_{2}^{\ast}(x,y)\leq C_{2}d_{1}^{\ast}(x,y)$. In a similar manner, $d_{1}^{\ast}\leq C_{1}d_{2}^{\ast}$. ∎
Example 5.9 (Euclidean distance on $\mathbb{R}^{n}$).
We denote by $d_{\mathbb{R}^{n}}$ the usual Euclidean distance on $\mathbb{R}^{n}$. It is well known and easy to verify that $d_{\mathbb{R}^{n}}^{\ast}=d_{\mathbb{R}^{n}}$. Identifying $\mathbb{C}^{n}$ with $\mathbb{R}^{2n}$, we obtain $d^{\ast}_{\mathbb{C}^{n}}=d_{\mathbb{C}^{n}}$. ∎
Example 5.10 (the intrinsic distance induced by the finite discrete distance).
Let $X$ be a set having more than one point and $d_{0}\colon X^{2}\to[0,+\infty)$ is defined by
$d_{0}(x,y)\coloneqq\begin{cases}0,&x=y;\\ 1,&x\neq y.\end{cases}$
Since $d_{0}$ induces the discrete topology, there are no continuous paths between different points of $X$, and $d_{0}^{\ast}(x,y)=+\infty$ if $x\neq y$. ∎
Example 5.11 (square-root distance on $\mathbb{R}$).
Consider the distance $d_{\operatorname{sqrt}}\colon\mathbb{R}^{2}\to[0,+\infty)$, given by $d_{\operatorname{sqrt}}(x,y)\coloneqq\sqrt{|x-y|}$. The topology induced by $d_{\operatorname{sqrt}}$ is the usual topology in $\mathbb{R}$, since the set of open balls with respect to the Euclidean distance on $\mathbb{R}$ is the same as the set of open balls with respect to the distance $d_{\operatorname{sqrt}}$.
Let $x,y\in\mathbb{R}$. We will show that if $x\neq y$, then $d_{\operatorname{sqrt}}^{\ast}(x,y)=+\infty$. Suppose $x<y$ (the case if $x>y$ is similar). Let $\gamma\in\Gamma_{d_{\operatorname{sqrt}}}(x,y)$. Notice that $\gamma$ is continuous in the usual sense. Let $m\in\mathbb{N}$. For every $j$ in $\{0,\ldots,m\}$, we define
$z_{m,j}\coloneqq x+\frac{j}{m}(y-x).$
By the intermediate value theorem, there is $t$ in $[0,1]$ such that $\gamma(t)=z_{m,j}$. Define
$t_{m,m}\coloneqq 1,$
$t_{m,j}\coloneqq\inf\,\bigl\{t\in[0,1]\colon\ \gamma(t)=z_{m,j}\bigr\}\qquad(j\in\{0,\ldots,m-1\}).$
Because $\gamma$ is continuous, $\gamma(t_{m,j})=z_{m,j}$. Furthermore, the intermediate value theorem implies
$t_{m,0}<t_{m,1}<\cdots<t_{m,m}.$
Using the partition $P_{m}\coloneqq\{t_{m,0},t_{m,1},\dots,t_{m,m}\}$, we establish the following lower bound for $L(d_{\operatorname{sqrt}},\gamma)$:
$L(d_{\operatorname{sqrt}},\gamma)$ $\geq S(d_{\operatorname{sqrt}},\gamma,P_{m})=\sum_{j=0}^{m-1}d_{\operatorname{sqrt}}(\gamma(t_{m,j}),\gamma(t_{m,j+1}))$
$=\sum_{j=0}^{m-1}\sqrt{z_{m,j+1}-z_{m,j}}=\sum_{j=0}^{m-1}\sqrt{\frac{1}{m}(y-x)}=\sqrt{m(y-x)}.$
Taking the supremum over $m$ in $\mathbb{N}$ shows that $L(d_{\operatorname{sqrt}},\gamma)=+\infty$. Since $\gamma$ was an arbitrary path between $x$ and $y$, it follows from Definition 5.6, that $d_{\operatorname{sqrt}}^{\ast}(x,y)=+\infty$ if $x\neq y$.
Let $d_{0}\colon\mathbb{R}^{2}\to[0,+\infty)$ is defined as in Example 5.10. Although $d_{0}$ and $d_{\operatorname{sqrt}}$ are not topologically equivalent, their corresponding intrinsic distances are, in fact, equal: $d_{0}^{\ast}=d_{\operatorname{sqrt}}^{\ast}$. ∎
In the following example we develop the idea from Example 5.11 in such a manner that the induced intrinsic distance is finite.
Example 5.12 (quadrant of the plane with the usual distance added with the sqrt-angular pseudodistance).
Let $X$ be the closed top right quadrant of the closed unit disk on the complex plane:
$X\coloneqq\bigl\{z\in\mathbb{C}\colon\ |z|\leq 1,\ \operatorname{Re}(z)\geq 0,\ \operatorname{Im}(z)\geq 0\bigr\}.$
Given $z$ in $X$, let $\arg(z)$ be the principal argument of $z$ belonging to $0,\pi/2]$. We define $d\colon X^{2}\to[0,+\infty)$ by
$d(z,w)\coloneqq\begin{cases}|z|+|w|,&zw=0;\\[4.30554pt] |z-w|,&zw\neq 0,\ \arg(z)=\arg(w);\\[4.30554pt] |z-w|+\sqrt{|\arg(z)-\arg(w)|},&zw\neq 0,\ \arg(z)\neq\arg(w);\end{cases}$
Furthermore, we define $\rho\colon X^{2}\to[0,+\infty)$ by
$\rho(z,w)\coloneqq\begin{cases}|z-w|,&zw\neq 0,\ \arg(z)=\arg(w);\\[4.30554pt] |z|+|w|,&\text{otherwise}.\end{cases}$
Both $d$ and $\rho$ are distances on $X$. The topology induced by $d$ coincides with the usual topology of the set $X$ induced by the Euclidean distance. The topology induced by $\rho$ is different. Indeed, if $z\neq 0$ and $0<\delta<\min\{|z|,1-|z|\}$, then the ball $B_{\rho}(z,\delta)$ is in fact the segment $\{\lambda z\colon\ |z|-\delta<\lambda<|z|+\delta\}$.
Given $z,w$ in $X$ such that $z\neq 0$, $w\neq 0$, and $\arg(z)\neq\arg(w)$, and a path $\gamma\in\Gamma_{d}(z,w)$ such that $0\notin\gamma([0,1])$, we apply the reasoning from Example 5.11 to $\arg\circ\gamma$ and conclude that $L(d,\gamma)=+\infty$.
Therefore, if $\eta$ is a path of finite length from $z$ to $w$, then it must pass through 0. Moreover, some subpaths of $\eta$ must be contained in the segments joining $z$ with 0 and 0 with $w$. Figure 4 shows a couple of paths: one has finite length with respect to $d$ and the other has infinite length.
It follows that $d^{\ast}=\rho$. Consequently, $\rho^{\ast}=\rho$. In this example, $d$ is not topologically equivalent to $\rho$, but they induce the same intrinsic distance.
[…]
Figure 4: The set $X$ from Example 5.12, two points belonging to $X$ ($z$ and $w$), a path of finite length between them (blue) and a path of infinite length (red), with respect to $d$.
Example 5.13 (pseudologarithmic distance on the halfline).
Let $X\coloneqq(0,+\infty)$. Define $\rho\colon X^{2}\to[0,+\infty)$,
$\rho(x,y)\coloneqq\frac{2|x-y|}{x+y}.$
We are going to prove that
$\rho^{\ast}(x,y)=|\log(x)-\log(y)|.$ (5.2)
Consider the case where $x<y$ (the case $x>y$ is similar).
1. Let $\xi(t)\coloneqq x+t(y-x)$ be the straight-line path between $x$ and $y$, and let $Q\in\mathcal{P}([0,1])$. For each $m$ in $\mathbb{N}$, we add to $Q$ the points $k/m$, where $k\in\{0,1,\ldots,m\}$, and obtain a refined partition $R_{m}=\{t_{m,0},t_{m,1},\ldots,t_{m,m}\}$. For each $k$ in $\{0,1,\ldots,m\}$, we define $z_{m,k}\coloneqq\xi(t_{m,k})$. This setup leads to an explicit formula for the polygonal length of $\xi$ associated with $R_{m}$:
$S(\rho,\xi,Q)$ $\leq S(\rho,\xi,R_{m})=\sum_{k=0}^{m-1}\rho(\xi(t_{m,k}),\xi(t_{m,k+1}))=\sum_{k=0}^{m-1}\rho(z_{m,k},z_{m,k+1})$
$=\sum_{k=0}^{m-1}\frac{2(z_{m,k+1}-z_{m,k})}{z_{m,k+1}+z_{m,k}}=\sum_{k=0}^{m-1}\frac{z_{m,k+1}-z_{m,k}}{\frac{z_{m,k+1}+z_{m,k}}{2}}.$
The resulting sum is a Riemannian sum for the integral of the function $f(t)=\frac{1}{t}$ over the interval $[x,y]$. The points $z_{m,0},z_{m,1},\ldots,z_{m,m}$ form a Riemannian partition of $[x,y]$, and its mesh tends to 0 as $m\to\infty$:
$\min_{k\in\{0,\ldots,m-1\}}\bigl(z_{m,k+1}-z_{m,k}\bigr)=(y-x)\,\min_{k\in\{0,\ldots,m-1\}}\bigl(t_{m,k+1}-t_{m,k}\bigr)\leq\frac{y-x}{m}.$
Therefore,
$S(\rho,\xi,Q)\leq\lim_{m\to\infty}S(\rho,\xi,R_{m})=\int_{x}^{y}\frac{\mathrm{d}{}t}{t}=\log(y)-\log(x).$
Since $Q$ was an arbitrary element of $\mathcal{P}([0,1])$, we have proven that $L(\rho,\xi)\leq\log(y)-\log(x)$.
2. For an arbitrary $\gamma$ in $\Gamma_{\rho}(x,y)$, we can construct a partition $P_{m}=\{t_{m,0},\ldots,t_{m,m}\}$ with $\gamma(t_{m,k})=x+k(y-x)/m$, as in Example 5.11, and obtain
$L(\rho,\gamma)\geq S(\rho,\gamma,P_{m})=\sum_{k=0}^{m-1}\rho\left(x+k\frac{y-x}{m},x+(k+1)\frac{y-x}{m}\right).$
The right-hand side is also a Riemannian sum for the integral $\int_{x}^{y}\mathrm{d}{}t/t$. Notice that the points $\gamma(t_{m,k})$, where $k\in\{0,\ldots,m\}$, form a Riemannian partition of $[x,y]$, and the mesh of this partition equals (yx)/m(y-x)/m. Passing to the limit as $m\to\infty$ we get
$L(\rho,\gamma)\geq\int_{x}^{y}\frac{\mathrm{d}{}t}{t}=\log(y)-\log(x).$
Thereby, we have proven (5.2). ∎
Example 5.14 (pseudologarithmic distance on a segment).
Let $X\coloneqq[1,2]$. Define $\rho\colon X^{2}\to[0,+\infty)$ as
$\rho(x,y)\coloneqq\frac{2|x-y|}{x+y}.$
In other words, we consider a metric subspace of the metric space from Example 5.13. It is easy to see that $\rho^{\ast}(x,y)=|\log(x)-\log(y)|$. Let $d_{[1,2]}$ be the Euclidean distance on $[1,2]$. It is easy to see that $\rho\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{[1,2]}$, but $\rho^{\ast}\neq d_{[1,2]}=d_{[1,2]}^{\ast}$. Moreover, there is no $C>0$ such that $\rho^{\ast}=Cd_{[1,2]}^{\ast}$. ∎
Example 5.15 (pseudohyperbolic distance on the halfplane).
Let $\mathbb{H}$ be the upper halfplane of the complex plane:
$\mathbb{H}\coloneqq\{z\in\mathbb{C}\colon\ \operatorname{Im}(z)>0\}.$
Define $\rho_{\mathbb{H}}\colon\mathbb{H}^{2}\to[0,+\infty)$,
$\rho_{\mathbb{H}}(z,w)\coloneqq\left|\frac{w-z}{w-\overline{z}}\right|.$ (5.3)
For numbers on the imaginary axis, $\rho_{\mathbb{H}}$ is the distance from Example 5.13, divided by 2:
$\rho_{\mathbb{H}}(ia,ib)=\frac{|a-b|}{a+b}\qquad(a,b>0).$
The distance $\rho_{\mathbb{H}}$ between two points is always greater or equal to the distance $\rho_{\mathbb{H}}$ between their projections to the imaginary axis:
$\rho_{\mathbb{H}}(z,w)\geq\rho_{\mathbb{H}}(i\,\operatorname{Im}(z),i\,\operatorname{Im}(w)).$ (5.4)
Indeed, for $z=x+iy$ and $w=u+iv$,
$\rho_{\mathbb{H}}(z,w)^{2}-\rho_{\mathbb{H}}(i\,\operatorname{Im}(z),i\,\operatorname{Im}(w))^{2}=\frac{(x-u)^{2}+(y-v)^{2}}{(x-u)^{2}+(y+v)^{2}}-\frac{(y-v)^{2}}{(y+v)^{2}}$
$\quad=\frac{(x-u)^{2}(y+v)^{2}-(x-u)^{2}(y-v)^{2}}{((x-u)^{2}+(y+v)^{2})(y+v)^{2}}=\frac{4vy(x-u)^{2}}{((x-u)^{2}+(y+v)^{2})(y+v)^{2}}\geq 0.$
Using (5.4) it is easy to see that the straight path between $ia$ and $ib$ is shorter than any other path. Therefore, by (5.2),
$\rho_{\mathbb{H}}^{\ast}(ia,ib)=\frac{1}{2}|\log(a)-\log(b)|\qquad(a,b>0).$
Given arbitrary $z,w$ in $\mathbb{H}$, it is possible to find a Möbius transform $\varphi$ of $\mathbb{H}$ and some numbers $a,b>0$ such that $\varphi(z)=ia$, $\varphi(w)=ib$; see, e.g., [2, Section 3.5]. Since $\rho_{\mathbb{H}}$ is invariant under the Möbius transforms of $\mathbb{H}$, $\rho_{\mathbb{H}}^{\ast}(z,w)=\rho_{\mathbb{H}}^{\ast}(ia,ib)$. After an explicit computation of $\varphi$, $a$, and $b$, one obtains
$\rho_{\mathbb{H}}^{\ast}(z,w)=\frac{1}{2}\log\frac{1+\rho_{\mathbb{H}}(z,w)}{1-\rho_{\mathbb{H}}(z,w)}=\frac{1}{2}\log\frac{|w-\overline{z}|+|w-z|}{|w-\overline{z}|-|w-z|}.$ (5.5)
Example 5.16 (pseudohyperbolic distance on the unit disk).
Let $\mathbb{D}\coloneqq\{z\in\mathbb{C}\colon\ |z|<1\}$. Define $\rho_{\mathbb{D}}\colon\mathbb{D}^{2}\to[0,+\infty)$,
$\rho_{\mathbb{D}}(z,w)\coloneqq\left|\frac{w-z}{1-w\overline{z}}\right|.$ (5.6)
A Cayley transform between $\mathbb{H}$ and $\mathbb{D}$ converts $\rho_{\mathbb{H}}$ into $\rho_{\mathbb{D}}$. Thereby, it is possible to show that
$\rho_{\mathbb{D}}^{\ast}(z,w)=\frac{1}{2}\log\frac{1+\rho_{\mathbb{D}}(z,w)}{1-\rho_{\mathbb{D}}(z,w)}=\frac{1}{2}\log\frac{|1-w\overline{z}|+|w-z|}{|1-w\overline{z}|-|w-z|}.$ (5.7)
It is well known (see, e.g., [19, Chapter 33] and [22, Section 1.5]) that $2\rho_{\mathbb{H}}^{\ast}$ is the geodesic distance associated with the Poincaré halfplane metric $\mathrm{d}{}x\,\mathrm{d}{}y/y^{2}$, and $2\rho_{\mathbb{D}}^{\ast}$ is the geodesic distance associated with the Poincaré disc metric $\mathrm{d}{}x\,\mathrm{d}{}y/(1-x^{2}-y^{2})^{2}$. Nevertheless, we emphasize that the intrinsic distances $\rho_{\mathbb{H}}^{\ast}$ and $\rho_{\mathbb{D}}^{\ast}$ in these classical examples can be computed without using Riemannian metrics.
6 Locally similar distances induce the same intrinsic distance
In this section, we prove the main result of the paper (Theorem 6.6).
We start with the following elementary lemma which provides a sufficient condition for one interval to be a subset of another. Figure 5 illustrates such intervals. Lemma 6.1 is useful in the proof of Lemma 6.2.
Lemma 6.1.
Let $x,y\in\mathbb{R}$ and $r,s>0$ such that one of the following conditions holds:
  1. (a)
    $y<x$ and $x+r<y+s$;
  2. (b)
    $x<y$ and ys<xry-s<x-r.
Then, $(x-r,x+r)\subseteq(y-s,y+s)$.
Proof.
We only consider the case (a). First, we compare the radii of the intervals:
s>(xy)+r>r.s>(x-y)+r>r.
Then, we make a conclusion about the left extremes: ys<xry-s<x-r. ∎
$\scriptstyle y-s$$\scriptstyle y+s$$\scriptstyle y$$\scriptstyle x-r$$\scriptstyle x+r$$\scriptstyle x$
$\scriptstyle y-s$$\scriptstyle y+s$$\scriptstyle y$$\scriptstyle x$$\scriptstyle x-r$$\scriptstyle x+r$
Figure 5: Lemma 6.1, case (a) (above) and (b) (below).
Next, we need the following fact about open covers and Riemannian partitions of $[a,b]$.
Lemma 6.2.
Let $\in\mathbb{R}, a<ba<b$, and let $R\colon[a,b]\to(0,+\infty)$. Then, there exist $m$ in $\mathbb{N}$ and $t_{0},\ldots,t_{m}$ in $[0,1]$ such that $a=t_{0}<t_{1}<\cdots<t_{m}=b$ and
$t_{k}-t_{k-1}<\max\{R(t_{k-1}),R(t_{k})\}\qquad(1\leq k\leq m).$ (6.1)
Proof.
Let $\mathcal{Q}$ be the subset of $\mathcal{P}([a,b])$ such that the corresponding $R(x)/2$-neighborhoods of its elements cover $[a,b]$:
$\mathcal{Q}\coloneqq\left\{P\in\mathcal{P}([a,b])\colon\quad[a,b]\subseteq\bigcup_{x\in P}\left(x-\frac{R(x)}{2},x+\frac{R(x)}{2}\right)\right\}.$
By the Heine–Borel theorem, $\mathcal{Q}\neq\emptyset$. Assigning to each set $P$ belonging to $\mathcal{Q}$ its cardinality #P\#P, we get a nonempty set $\{\#P\colon\ P\in\mathcal{Q}\}$ of natural numbers. By the induction principle, it has a minimal element. We denote it by $m+1$:
$m\coloneqq\min\bigl(\{\#P\colon\ P\in\mathcal{Q}\}\bigr)-1.$
Then, we choose in $\mathcal{Q}$ an element $Q$ of minimal size $m+1$.
Let $Q=\{t_{0},t_{1},\ldots,t_{m}\}$, where $a=t_{0}<t_{1}<\ldots<t_{m}=b$. For brevity, we set $r_{k}\coloneqq R(t_{k})/2$ for each $k$. We claim that for every $k$ in $\{1,\ldots,m\}$,
$t_{k-1}+r_{k-1}>t_{k}-r_{k}.$ (6.2)
In other words, we consider the list of intervals $((t_{k}-r_{k},t_{k}+r_{k}))_{k=0}^{m}$, and we are going to prove that each pair of consecutive intervals from this list intersect.
Reasoning by reductio ad absurdum we suppose that $k$ is an element of $\{1,\ldots,m\}$ such that
$t_{k-1}+r_{k-1}\leq t_{k}-r_{k}.$ (6.3)
Since $t_{k}-r_{k}\in[a,b]$, the following inequalities hold:
$a\leq t_{k-1}<t_{k-1}+r_{k-1}\leq t_{k}-r_{k}<t_{k}\leq b.$
Since $Q\in\mathcal{Q}$, the point $t_{k}-r_{k}$ must be covered by some of the intervals $\left(x-\frac{R(x)}{2},x+\frac{R(x)}{2}\right)$, where $x\in Q$. We choose $j$ in $\{0,1,\ldots,m\}$ such that
$t_{j}-r_{j}<t_{k}-r_{k}<t_{j}+r_{j}.$ (6.4)
Obviously, $j\neq k$. Moreover, due to (6.3), $j\neq k-1$.
First, consider the case where $j>k$. Then, $t_{j}>t_{k}$ and $t_{j}-r_{j}<t_{k}-r_{k}$. By Lemma 6.1,
$(t_{k}-r_{k},t_{k}+r_{k})\subseteq(t_{j}-r_{j},t_{j}+r_{j}).$
Notice that $k\in\{1,\ldots,m-1\}$ because $1\leq k<j\leq m$. Therefore, we can exclude $\{t_{k}\}$ from $Q$ without losing the covering property nor the property that $a,b\in Q$. So, $Q\setminus\{t_{k}\}$ belongs to $\mathcal{Q}$ and has fewer elements than $Q$, which contradicts the minimality of $Q$.
Now, consider the case where j<k1j<k-1. In this case, $t_{j}<t_{k-1}$. On the other hand, by (6.4) and (6.3),
$t_{k-1}+r_{k-1}\leq t_{k}-r_{k}<t_{j}+r_{j},$
By Lemma 6.1,
$(t_{k-1}-r_{k-1},t_{k-1}+r_{k-1})\subseteq(t_{j}-r_{j},t_{j}+r_{j}).$
Notice that $k-1\in\{1,\ldots,m-1\}$ because $0\leq j<k-1<k\leq m$. Excluding $t_{k-1}$ from $Q$ we get an element of $\mathcal{Q}$ with a smaller number of elements, which again contradicts the minimality of $Q$.
Therefore, (6.2) is true, and (6.2) easily implies (6.1):
$t_{k}-t_{k-1}<r_{k-1}+r_{k}=\frac{R(t_{k-1})+R(t_{k})}{2}\leq\max\{R(t_{k-1}),R(t_{k})\}.\qed$
Remark 6.3.
The property (6.1) from Lemma 6.2 has a simple geometrical meaning: for each $k$, at least one of the points $t_{k-1}$ and $t_{k}$ belongs to the interval associated with the other point. In other words, for each $k$ in $\{1,\ldots,m\}$, at least one of the following conditions is fulfilled:
  • (a)
    $t_{k}\in(t_{k-1}-R(t_{k-1}),t_{k-1}+R(t_{k-1}))$;
  • (b)
    $t_{k-1}\in(t_{k}-R(t_{k}),t_{k}+R(t_{k}))$.
Indeed, if $\max\{R(t_{k-1}),R(t_{k})\}=R(t_{k-1})$, then condition (a) holds. Otherwise, we get $\max\{R(t_{k-1}),R(t_{k})\}=R(t_{k})$, and (b) holds.
Remark 6.4.
The axiom of choice is unnecessary in this paper, and we avoid using it. It is well known that the Heine–Borel theorem can be proved without using the axiom of choice; see, e.g., Lebesgue’s proof explained in [16, Chapter 2, Section 5, Theorem 15]. In the proof of the forthcoming Lemma 6.5, to define a function $R\colon[a,b]\to(0,+\infty)$ without using the axiom of choice, we select $R(x)$ as the maximal number of the form $1/q$, where $q\in\mathbb{N}$, that has the needed property.
Lemma 6.5.
Let $X$ be a set and $d_{1},d_{2}$ be distances on $X$ such that $d_{1}\cong d_{2}$. Let $\in\mathbb{R}, a<ba<b$, and $\eta\in C_{d_{1}}([a,b],X)$. Then, for every $\varepsilon>0$, there exist $m$ in $\mathbb{N}$ and $t_{0},t_{1},\ldots,t_{m}$ in $[a,b]$ such that $a=t_{0}<t_{1}<\cdots<t_{m}=b$ and
$\sum_{k=1}^{m}d_{2}(\eta(t_{k-1}),\eta(t_{k}))\leq(1+\varepsilon)\sum_{k=1}^{m}d_{1}(\eta(t_{k-1}),\eta(t_{k})).$ (6.5)
Proof.
For each $t$ in $[a,b]$, by the definition of $\cong$, there exists an open neighborhood $W$ of $\eta(t)$ such that
$\forall u\in W\qquad d_{2}(u,\eta(t))<(1+\varepsilon)d_{1}(u,\eta(t)).$
Furthermore, by the continuity of $\eta$, there exists $r>0$ such that if $s\in[a,b]$ and |st|<r|s-t|<r, then $\eta(s)\in W$. Moreover, by the Archimedes–Eudoxus principle, there exists $q$ in $\mathbb{N}$ such that $1/q<r$. Thus, for each $t$ in $[a,b]$, the following set is nonempty:
$D(t)\coloneqq\Bigl\{q\in\mathbb{N}\colon\ \forall s\in[a,b]\cap(t-1/q,t+1/q)\quad d_{2}(\eta(s),\eta(t))<(1+\varepsilon)d_{1}(\eta(s),\eta(t))\Bigr\}.$
By the induction principle, $D(t)$ has a minimum element. We define $R\colon[a,b]\to(0,+\infty)$ by
$R(t)\coloneqq\frac{1}{\min(D(t))}.$
The definition of $R(t)$ implies that for each $t$ in $[a,b]$.
$\forall s\in[a,b]\cap(t-R(t),t+R(t))\qquad d_{2}(\eta(s),\eta(t))<(1+\varepsilon)d_{1}(\eta(s),\eta(t)).$ (6.6)
Applying Lemma 6.2, we find $m$ in $\mathbb{N}$ and $t_{0},\ldots,t_{m}$ in $[a,b]$ satisfying (6.1).
For each $k$ in $\{1,\ldots,m\}$, due to (6.1), we have at least one of the cases (a) or (b) from Remark 6.3.
In the case (a), we use (6.6) with $t_{k-1}$ instead of $t$ and $t_{k}$ instead of $s$, and we get
$d_{2}(\eta(t_{k-1}),\eta(t_{k}))<(1+\varepsilon)d_{1}(\eta(t_{k-1}),\eta(t_{k})).$ (6.7)
In the case (b), we apply (6.6) with $t_{k}$ instead of $t$ and $t_{k-1}$ instead of $s$, and we also get (6.7).
Finally, summing (6.7) over all $k$ in $\{1,\ldots,m\}$, we obtain (6.5). ∎
Theorem 6.6.
If $X$ is a set and $d_{1},d_{2}\colon X\to[0,+\infty)$ are distances on $X$ such that $d_{1}\cong d_{2}$, then $d_{1}^{\ast}=d_{2}^{\ast}$.
Proof.
We suppose that $d_{1},d_{2}\colon X\to[0,+\infty)$ such that $d_{1}\cong d_{2}$. Let $x,y\in X$. Since $d_{1}$ and $d_{2}$ are topologically equivalent, $\Gamma_{d_{1}}(x,y)=\Gamma_{d_{2}}(x,y)$. Let $\gamma\in\Gamma_{d_{1}}(x,y)$ and $\varepsilon>0$. Given a partition $P$ in $\mathcal{P}$, we enumerate its elements in the ascending order: $P=\{p_{0},p_{1},\ldots,p_{N}\}$, $0=p_{0}<p_{1}<\cdots<p_{N}=1$. Then, for each $j$ in $\{1,\ldots,N\}$, we apply Lemma 6.5 to the subpath $\eta_{j}\coloneqq\gamma|_{[p_{j-1},p_{j}]}$. Thereby, we get $m_{j}$ in $\mathbb{N}$ and a list of real numbers $t_{j,0},t_{j,1},\dots,t_{j,m_{j}}$ such that
$p_{j-1}=t_{j,0}<t_{j,1}<\cdots<t_{j,m_{j}}=p_{j}$
and
$\sum_{k=1}^{m_{j}}d_{2}(\gamma(t_{j,k-1}),\gamma(t_{j,k}))\leq(1+\varepsilon)\sum_{k=1}^{m_{j}}d_{1}(\gamma(t_{j,k-1}),\gamma(t_{j,k})).$ (6.8)
Let $Q$ be the partition consisting of all these points $t_{j,k}$, where $1\leq j\leq N$ and $1\leq k\leq m_{j}$. Then, $P\subseteq Q$ and $Q\in\mathcal{P}$. Summing (6.8) over $j$ in $\{1,\ldots,N\}$ we obtain
$S(\gamma,Q,d_{2})\leq(1+\varepsilon)S(\gamma,Q,d_{1}).$
Therefore,
$S(\gamma,P,d_{2})\leq S(\gamma,Q,d_{2})\leq(1+\varepsilon)S(\gamma,Q,d_{1})\leq(1+\varepsilon)L(d_{1},\gamma).$
We conclude that $L(d_{2},\gamma)\leq L(d_{1},\gamma)$. Taking the infimum over $\gamma$ in $\Gamma_{d_{1}}(x,y)$, we get $d_{2}^{\ast}(x,y)\leq d_{1}^{\ast}(x,y)$. The inequality $d_{1}^{\ast}(x,y)\leq d_{2}^{\ast}(x,y)$ follows with a similar reasoning, swapping the roles of $d_{1}$ and $d_{2}$ in Lemma 6.5. ∎
Corollary 6.7.
If $X$ is a set and $d_{1},d_{2}$ are distances on $X$ such that $d_{1}\cong d_{2}$ and $d_{2}^{\ast}=d_{2}$, then $d_{1}^{\ast}=d_{2}$.
In Examples 5.13 and 5.16, the distances have the property $d^{\ast}\cong d$.
Problem 6.8.
Find a necessary and sufficient condition for $d^{\ast}\cong d$.
Notice that $d_{1}\cong d_{2}$ is not a necessary condition for $d_{1}^{\ast}=d_{2}^{\ast}$. We have seen it in Examples 5.11 and 5.12. In that examples, the initial distances where not topologically equivalent.
In the following Example 6.9, two distances $d$ and $\rho$ induce the same topology and the same finite intrinsic distance $d^{\ast}=\rho^{\ast}$, but $d$ and $\rho$ are not locally similar.
Example 6.9.
Let $X$ be the following “comb” in $\mathbb{R}^{2}$; see Figure 6:
$X\coloneqq([0,1]\times\{0\})\cup\left(\bigcup_{q\in\mathbb{N}}\left\{\frac{1}{q}\right\}\times[0,1]\right).$
Notice that the ordinate axis $\{0\}\times\mathbb{R}$ only intersects $X$ at the origin.
Let $d$ be the Euclidean distance on $X$, inherited from $\mathbb{R}^{2}$. Define $\rho\colon X^{2}\to[0,+\infty)$,
$\rho(a,b)\coloneqq\begin{cases}|a_{2}-b_{2}|,\quad\text{if}\quad a_{1}=b_{1};\\ |a_{1}-b_{1}|+a_{2}+b_{2},\quad\text{if}\quad a_{1}\neq b_{1};\end{cases}\qquad(a=(a_{1},a_{2}),b=(b_{1},b_{2})\in X).$
It is easy to see that $\rho$ is a distance on $X$. The distances $d$ and $\rho$ are topologically equivalent; this can be proved by comparing the balls $B_{d}(a,\delta_{1})$ and $B_{\rho}(a,\delta_{2})$ for each $a$ in $X$, with small $\delta_{1}$ and $\delta_{2}$.
Furthermore, $d^{\ast}=\rho$. In other words, we claim that $d^{\ast}(a,b)=\rho(a,b)$ for every $a$, $b$ in $X$. For brevity, we illustrate the idea of the proof for the particular case where $a=(0,0)$ and $b=(1/q,y)$, $q\in\mathbb{N}$, $0<y\leq 1$. If $\gamma\colon[0,1]\to X$ is a continuous path joining $a$ with $b$, then, by connectivity arguments, there exists $u$ in $(0,1)$ such that $\gamma(u)=(1/q,0)$. It is easy to see that $L(d,\gamma|_{[0,u]})\geq 1/q$ and $L(d,\gamma|_{[u,1]})\geq y$; here the domains of the paths are $[0,u]$ and $[u,1]$; see Remark 5.7. Therefore, $L(d,\gamma)\geq\rho(a,b)$ for every $\gamma$ in $\Gamma_{d}(a,b)$, and it is easy to conclude that $d^{\ast}(a,b)=\rho(a,b)$. In particular, $d\leq\rho$.
Let us show that $\rho$ is not locally similar to $d$. Indeed, taking $a=(0,0)$, $b_{q}=(1/q,1/q)$, and $c_{q}=(1/q,0)$ for $q$ in $\mathbb{N}$, we get
$\frac{\rho(a,b_{q})}{d(a,b_{q})}=\frac{2/q}{\sqrt{2}/q}=\sqrt{2},\qquad\frac{\rho(a,c_{q})}{d(a,c_{q})}=\frac{1/q}{1/q}=1.$
Taking $q$ large enough, we can get $b_{q}$ and $c_{q}$ arbitrarily close to $a$. Thus, the quotient $\rho(a,x)/d(a,x)$ does not have any limit as $x$ tends to $a$. More precisely, the superior and inferior limits defined in (2.7) are $L_{\operatorname{id}_{X}}(a)=\sqrt{2}$ and $\ell_{\operatorname{id}_{X}}(a)=1$. This means that $\operatorname{id}_{X}$ has distortion $\sqrt{2}$ at the point $a$. ∎
$\scriptstyle(0,0)$$\scriptstyle(1,0)$$\scriptstyle\left(\frac{1}{2},0\right)$$\scriptstyle\left(\frac{1}{2},\frac{1}{2}\right)$$\scriptstyle\left(\frac{1}{3},0\right)$
Figure 6: Part of the comb from Example 6.9. The light-green thick lines illustrate the shortest path from $(0,0)$ to $\left(\frac{1}{2},\frac{1}{2}\right)$.
7 Examples related to kernels
Let $X$ be a set. A function $K\colon X^{2}\to\mathbb{C}$ is called a kernel on $X$ if for every $m$ in $\mathbb{N}$, every $x_{1},\ldots,x_{m}$ in $X$ and every $\lambda_{1},\ldots,\lambda_{m}$ in $\mathbb{C}$,
$\sum_{r,s=1}^{m}\overline{\lambda_{r}}\,\lambda_{s}\,K(x_{r},x_{s})\geq 0.$ (7.1)
By the Moore–Aronszajn theorem (see [4, Part 1, Section 4] or [15, Theorem 2.14]), every kernel corresponds to a uniquely determined reproducing kernel Hilbert space (RKHS). Conversely, every RKHS has a uniquely defined kernel. So, speaking about kernels is equivalent to speaking about RKHS.
A function $K\colon X^{2}\to\mathbb{C}$ is called a strictly positive kernel on $X$ if for every $m$ in $\mathbb{N}$, every list of pairwise different points $x_{1},\ldots,x_{m}$ in $X$ and every $\lambda_{1},\ldots,\lambda_{m}$ in $\mathbb{C}$, not all zero,
$\sum_{r,s=1}^{m}\overline{\lambda_{r}}\,\lambda_{s}\,K(x_{r},x_{s})>0.$ (7.2)
See various characterizations of strictly positive kernels in [15, Section 3.2].
Agler and McCarthy [1, Lemma 9.9] proved that the following function $d_{K}$ is a distance on $X$. See also Arcozzi, Rochberg, Sawyer, and Wick [3].
Definition 7.1 (the distance associated with a kernel).
Let $X$ be a set and $K$ be a strictly positive kernel on $X$. Define $d_{K}\colon X^{2}\to[0,+\infty)$ by
$d_{K}(x,y)\coloneqq\sqrt{1-\frac{|K(x,y)|^{2}}{K(x,x)K(y,y)}}.$ (7.3)
In this section, we apply the previous tools, including Theorem 6.6 and Corollary 6.7, to compute $d_{K}^{\ast}$ for some examples.
We start with two classical examples (Examples 7.2 and 7.3) that are usually studied with the help of the Bergman metrics (for these two examples, the Bergman metrics are constant multiples of the Poincaré metrics), see [6, 22]. We show that these examples can be treated with more elementary tools.
Example 7.2 (Bergman space on the upper halfplane).
Consider the Bergman space of the analytic square-integrable functions on the upper halfplane with the Lebesgue measure. The reproducing kernel of this space is
$K(w,z)=\frac{1}{\pi(w-\overline{z})^{2}}.$
The associated distance is
$d_{K}(w,z)=\left|\frac{w-z}{w-\overline{z}}\right|\sqrt{2-\left|\frac{w-z}{w-\overline{z}}\right|^{2}}.$
So, $d_{K}$ can be written as the composition $f\circ\rho_{\mathbb{H}}$, where $\rho_{\mathbb{H}}$ is as defined in (5.3), and
$f(t)\coloneqq t\sqrt{2-t^{2}}.$
It is easy to see that $f$ satisfies conditions of Corollary 4.3 with $Q=\sqrt{2}$. Therefore, $d_{K}\cong\sqrt{2}\,\rho_{\mathbb{H}}$. By Theorem 6.6, $d_{K}^{\ast}=\sqrt{2}\,\rho_{\mathbb{H}}^{\ast}$. Recall that $\rho_{\mathbb{H}}^{\ast}$ is the hyperbolic distance on $\mathbb{H}$, see Example 5.15. ∎
Example 7.3 (Bergman space on the unit disk).
Consider the Bergman space of the analytic square-integrable functions on $\mathbb{D}$ with the Lebesgue measure; see, e.g., [6, 22]. The reproducing kernel of this space is
$K(w,z)=\frac{1}{\pi(1-w\overline{z})^{2}}.$
The associated distance is
$d_{K}(z,w)=\left|\frac{z-w}{1-\overline{z}w}\right|\sqrt{2-\left|\frac{z-w}{1-\overline{z}w}\right|^{2}}.$
Note that $d_{K}=f\circ\rho_{\mathbb{D}}$, where $\rho_{\mathbb{D}}$ is as in (5.6) and
$f(t)\coloneqq t\sqrt{2-t^{2}}.$
Reasoning as in Example 7.2, by Corollary 4.3, $d_{K}\cong\sqrt{2}\,\rho_{\mathbb{D}}$. By Theorem 6.6, $d_{K}^{\ast}=\sqrt{2}\,\rho^{\ast}_{\mathbb{D}}$. ∎
Example 7.4 (Hardy space and Szegő kernel on the disk).
This example is studied in [3]. The reproducing kernel of the Hardy space, known as Szegő kernel, is
$K(w,z)=\frac{1}{1-w\overline{z}}.$
The associated distance is the pseudohyperbolic distance: $d_{K}=\rho_{\mathbb{D}}$. Therefore, $d_{K}^{\ast}=\rho_{\mathbb{D}}^{\ast}$. ∎
Example 7.5 (the Gaussian kernel).
Let $\sigma>0$. We consider the following kernel on $\mathbb{R}^{n}$:
$K(x,y)=\exp(-\sigma^{2}\|x-y\|^{2}).$
The corresponding RKHS was described by Saitoh [17] and by Steinwart, Hush, and Scovel [18]. In this example,
$d_{K}(x,y)=\sqrt{1-\exp(-2\sigma^{2}\|x-y\|^{2})}.$
So, $d_{K}$ can be written as the composition $f\circ d_{\mathbb{R}^{n}}$, where
$f(t)\coloneqq\sqrt{1-\operatorname{e}^{-2\sigma^{2}t^{2}}}.$ (7.4)
It is easy to see that $f$ satisfies the conditions from Corollary 4.3 with $Q=\sqrt{2}\sigma$. Therefore, $d_{K}\cong\sqrt{2}\sigma d_{\mathbb{R}^{n}}$. By Theorem 6.6 and Example 5.9, $d_{K}^{\ast}=\sqrt{2}\sigma d_{\mathbb{R}^{n}}^{\ast}=\sqrt{2}\sigma d_{\mathbb{R}^{n}}$. ∎
Example 7.6 (Bargmann–Segal–Fock space on the complex plane).
For the definition and properties of this space, see [21]. Its reproducing kernel is
$K(w,z)=\operatorname{e}^{w\overline{z}}.$
In this example, $d_{K}\colon\mathbb{C}^{2}\to[0,+\infty)$ is defined by
$d_{K}(w,z)=\sqrt{1-\operatorname{e}^{-|z-w|^{2}}}\qquad(w,z\in\mathbb{C}).$
We notice that $d_{K}=f\circ d_{\mathbb{C}}$, where $d_{\mathbb{C}}$ is the Euclidean distance on $\mathbb{C}$ and
$f(t)\coloneqq\sqrt{1-\operatorname{e}^{-t^{2}}}.$
By Corollary 4.3, $d_{K}\cong d_{\mathbb{C}}$. By Theorem 6.6 and Example 5.9, $d_{K}^{\ast}=d_{\mathbb{C}}^{\ast}=d_{\mathbb{C}}$. ∎
Example 7.7 (polyanalytic Bargmann–Segal–Fock space on the complex plane).
This is a generalization of Example 7.6. Let $m\geq 1$. Consider the space of the $m$-analytic functions on $\mathbb{C}$, square integrable with respect to the Gaussian measure. The corresponding reproducing kernel was computed by Askour, Intissar, and Mouayn [5]:
$K(w,z)=\operatorname{e}^{w\,\overline{z}}L^{1}_{m-1}(|w-z|^{2}).$ (7.5)
See also Youssfi [20] for another proof of (7.5), in the multidimensional case. Here $L^{1}_{m-1}$ denotes the associated Laguerre(-Sonine) polynomial of degree m1m-1, with parameter $1$:
$L^{1}_{k}(x)=\sum_{j=0}^{k}(-1)^{j}\binom{k+1}{k-j}\frac{x^{j}}{j!}.$ (7.6)
The associated distance is
$d_{K}(w,z)=\sqrt{1-\frac{1}{m^{2}}\operatorname{e}^{-|w-z|^{2}}L^{1}_{m-1}(|w-z|^{2})^{2}}\qquad(w,z\in\mathbb{C}).$ (7.7)
We see that $d_{K}=f\circ d_{\mathbb{C}}$, where
$f(t)\coloneqq\sqrt{1-\frac{1}{m^{2}}\operatorname{e}^{-t^{2}}L^{1}_{m-1}(t^{2})^{2}}.$
It follows from (7.6) that $L_{m-1}^{1}(t^{2})^{2}=m^{2}-m^{2}(m-1)t^{2}+O(t^{4})$. Therefore, an elementary analysis shows that $f$ satisfies conditions of Proposition 4.2, with $Q=\sqrt{m}$. Therefore, $d_{K}\cong\sqrt{m}\,d_{\mathbb{C}}$. By Theorem 6.6 and Example 5.9, $d_{K}^{\ast}=\sqrt{m}\,d_{\mathbb{C}}^{\ast}=\sqrt{m}\,d_{\mathbb{C}}$. ∎
Example 7.8 (the space of band-limited functions).
Given $f$ in $L^{2}(\mathbb{R})$, we denote by $\widehat{f}$ the Fourier–Plancherel transform of $f$, with the factor $\exp(-2\pi\operatorname{i}xt)$ under the integral. Let $A>0$. The space of band-limited functions with halfwidth $A$, also known as the Paley–Wiener space of halfwidth $A$, is defined by
$\mathcal{PW}_{A}\coloneqq\bigl\{\widehat{g}\colon\ g\in L^{2}(\mathbb{R}),\ \mu(\{x\in\mathbb{R}\setminus[-A,A]\colon\ g(x)\neq 0\})=0\bigr\}.$
This space is considered with the inner product $\langle\widehat{g_{1}},\widehat{g_{2}}\rangle_{\mathcal{PW}_{A}}\coloneqq\langle g_{1},g_{2}\rangle_{L^{2}(\mathbb{R})}$ which coincides with $\langle\widehat{g_{1}},\widehat{g_{2}}\rangle_{L^{2}(\mathbb{R})}$ by the Plancherel theorem. We denote by $\operatorname{sinc}$ the normalized cardinal sine function defined on $\mathbb{R}$ by
$\operatorname{sinc}(x)\coloneqq\begin{cases}1,&x=0;\\ \frac{\sin(\pi x)}{\pi x},&x\in\mathbb{R}\setminus\{0\}.\end{cases}$
Using the formula
$(\widehat{1_{[-A,A]}})(t)=2A\operatorname{sinc}(2At)\qquad(t\in\mathbb{R})$ (7.8)
it is easy to see (e.g., [14, Section 4.3, Example 4.2], [15, Subsection 1.3.2]) that $\mathcal{PW}_{A}$ is a RKHS on $\mathbb{R}$, and its reproducing kernel is
$K(s,t)=2A\operatorname{sinc}(2A(s-t)).$ (7.9)
The associated distance is
$d_{K}(s,t)=\sqrt{1-\operatorname{sinc}^{2}(2A(t-s))}.$ (7.10)
Notice that $d_{K}=f\circ\left(\frac{2A\pi}{\sqrt{3}}\,d_{\mathbb{R}}\right)$, where
$f(t)\coloneqq\sqrt{1-\operatorname{sinc}^{2}\frac{\sqrt{3}\,t}{\pi}}.$
A simple calculation with power series yields $f^{\prime}(0)=1$. Using the identity $\sin^{2}(x)=\frac{1-\cos(2x)}{2}$, we get
$\operatorname{sinc}^{2}(x)=\frac{1-\cos(2\pi x)}{2(\pi x)^{2}}\leq 1-\frac{(\pi x)^{2}}{3}+\frac{2(\pi x)^{4}}{45}.$
If $t\in[0,1]$, then
$f(t)=\sqrt{1-\operatorname{sinc}^{2}\frac{\sqrt{3}\,t}{\pi}}\geq\sqrt{t^{2}-\frac{2t^{4}}{5}}=\sqrt{t^{2}\left(1-\frac{2t^{2}}{5}\right)}\geq\frac{t}{2}.$
Recall that $|\operatorname{sinc}(x)|\leq\frac{1}{\pi x}$. If $t>1$, then
$f(t)=\sqrt{1-\operatorname{sinc}^{2}\frac{\sqrt{3}\,t}{\pi}}\geq\sqrt{1-\frac{1}{3t^{2}}}>\frac{1}{2}.$
Joining the two cases, we get
$f(t)\geq\frac{1}{2}\min\{t,1\}\qquad(t>0).$
This means that $f$ satisfies the hypothesis of Proposition 4.2 with $Q=1$. So, $d_{K}\cong\frac{2A\pi}{\sqrt{3}}d_{\mathbb{R}}$, and $d_{K}^{\ast}=\frac{2A\pi}{\sqrt{3}}d_{\mathbb{R}}^{\ast}=\frac{2A\pi}{\sqrt{3}}d_{\mathbb{R}}$. ∎
Example 7.9 (a Sobolev space).
We denote by $\operatorname{AC}([0,1])$ the set of all absolutely continuous complex-valued functions on $[0,1]$. Consider the vector space
$\mathcal{H}\coloneqq\bigl\{f\in\operatorname{AC}([0,1])\colon\ f(0)=0,\ f(1)=0,\ f^{\prime}\in L^{2}([0,1])\bigr\},$
with the inner product $\langle f,g\rangle_{\mathcal{H}}\coloneqq\langle f^{\prime},g^{\prime}\rangle_{L^{2}([0,1])}$. It is known [15, Subsection 1.3.1] that $\mathcal{H}$ is an RKHS, and its reproducing kernel is
$K(x,y)=\begin{cases}(1-x)y,&y\leq x;\\ (1-y)x,&x<y.\end{cases}$ (7.11)
The distance associated to $K$ is
$d_{K}(x,y)=\begin{cases}0,&x=y;\\ \sqrt{\frac{y-x}{(1-y)x}},&y>x;\\ \sqrt{\frac{x-y}{(1-x)y}},&x>y.\end{cases}$ (7.12)
Since $x\leq 1$ and $y\leq 1$,
$d_{K}(x,y)\geq\sqrt{|x-y|}.$
Reasoning similarly to Example 5.11, we see that
$d_{K}^{\ast}(x,y)=\begin{cases}0,&x=y;\\ +\infty,&x\neq y.\end{cases}$ (7.13)
Example 7.10 (the min-kernel on the unit interval).
Consider the vector space
$\mathcal{H}\coloneqq\bigl\{f\in\operatorname{AC}([0,1])\colon\ f(0)=0,\ f^{\prime}\in L^{2}([0,1])\bigr\}$
with the inner product $\langle f,g\rangle_{\mathcal{H}}\coloneqq\langle f^{\prime},g^{\prime}\rangle_{L^{2}([0,1])}$. It is known [14, Section 4.3, Example 4.3] that $\mathcal{H}$ is an RKHS, and its reproducing kernel is
$K(x,y)=\min\{x,y\}.$ (7.14)
This kernel corresponds to the covariance for the Wiener process. The induced distance is
$d_{K}(x,y)=\begin{cases}0,&x=y;\\ \sqrt{\frac{y-x}{y}},&y>x;\\ \sqrt{\frac{x-y}{x}},&x>y.\end{cases}$ (7.15)
Since $x\leq 1$ and $y\leq 1$,
$d_{K}(x,y)\geq\sqrt{|x-y|}.$
Similarly to Example 7.9, $d_{K}^{\ast}$ is given by (7.13). ∎
Acknowledgements
We are grateful to Prof. John Harvey (https://orcid.org/0000-0001-9211-0060) for the bibliographic reference [13] which we use in Section 3, Definition 3.6. We are also grateful to Prof. David José Fernández-Bretón (https://orcid.org/0000-0001-8084-055X) for explaining to us the ideas from Remark 6.4. We are grateful to Prof. Miguel Angel Rodriguez Rodriguez (https://orcid.org/0000-0002-5124-8271) for revising a draft of this paper and correcting some minor errors. We are grateful to Omer Cantor (https://orcid.org/0009-0007-2969-7170) for Example 3.11.
Funding
The second author has been supported by Proyecto SECIHTI (Mexico) “Ciencia de Frontera” FORDECYT-PRONACES/61517/2020 and by IPN-SIP projects (Instituto Politécnico Nacional, Mexico). The third author has been supported by SECIHTI (Mexico) scholarship.